OpenStudy (anonymous):

If sec theta+tan theta = P. PT sin theta=P^2-1/P^2+1

OpenStudy (anonymous):

first of fall sec theta=1/cos theta and tan thata= sintheta/costheta than you pu this value in rhs from that you can proved that

OpenStudy (anonymous):

u got it ?

OpenStudy (jhannybean):

No i don't understand what the last part means..

OpenStudy (jhannybean):

\[\large \sec(\theta)+ \tan(\theta) = P\] so \[\sin(\theta) = \frac{P^2-1}{P^2+1} \]

OpenStudy (jhannybean):

Thats the question?

hartnn (hartnn):

lol., who's question is this?

OpenStudy (anonymous):

ohk 1/cos + sin/cos=p then you take the lcm 1+ sin/cos is the value of p then you put this value in p2-1/p2+1

OpenStudy (jhannybean):

Okay sonot trying to evaluate RHS but LHs....

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

but lhs..........

OpenStudy (anonymous):

hey r u bsy

OpenStudy (jhannybean):

\[\sec(\theta) + \tan(\theta)\]\[\frac{1}{\cos(\theta)} +\frac{\sin(\theta)}{\cos(\theta)}\]\[\frac{1+\sin(\theta)}{\cos(\theta)}\] evaluating left hand side...but now what?

OpenStudy (jhannybean):

You can square the top and bottom, right?

OpenStudy (anonymous):

u put this value in place of p p2-1/p2+1

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ohk bye bye see you later

OpenStudy (jhannybean):

:|

OpenStudy (anonymous):

hey friends bye bye

OpenStudy (anonymous):

anyone have any doubt

OpenStudy (anonymous):

i will be back soon

OpenStudy (callisto):

\[\frac{P^2-1}{P^2+1}\]\[=\frac{(sec\theta+tan\theta)^2-1}{(sec\theta+tan\theta)^2+1}\]\[=\frac{sec^2\theta+2sec\theta+tan^2\theta-1}{sec^2\theta+2sec\theta+tan^2\theta+1}\]\[=\frac{2sec\theta tan\theta+2tan^2\theta}{2sec^2\theta+2sec\theta tan\theta}\]\[=\frac{2tan\theta(sec\theta +tan\theta)}{2sec\theta(sec\theta+ tan\theta)}\]\[=sin\theta\]Can I use Q.E.D. here? :D

OpenStudy (jhannybean):

ohIsee now lol.

OpenStudy (jhannybean):

I got it xD

OpenStudy (callisto):

Sorry for the typos in the third line, it should be \[=\frac{sec^2θ+2secθtanθ+tan^2θ−1}{sec^2θ+2secθtanθ+tan^2θ+1}\] It can be re arranged as \[=\frac{2secθtanθ+tan^2θ+(sec^2θ−1)}{sec^2θ+2secθtanθ+(tan^2θ+1)}\] By \(sec^2x=tan^2x+1\), which can be derived from \(sin^2x+cos^2x=1\) \[=\frac{2secθtanθ+tan^2θ+tan^2x}{sec^2θ+2secθtanθ+sec^2θ}\]

OpenStudy (callisto):

Ughhh Another typo on the last line. It should be \[=\frac{2secθtanθ+tan^2θ+tan^2\theta}{sec^2θ+2secθtanθ+sec^2θ}\]

OpenStudy (jhannybean):

lol its ok :)